Calculate the wavelength of light emitted when each of the following transitions occur in the hydrogen atom.



n = 5 to n = 3



_____________ nm

(Answer should be reported with 4 sig figs)
1
Expert's answer
2020-11-18T13:44:24-0500

λ-1= R(n-2final-n-2initial)

n final=3 while n initial=5

R=1.097*107m-1

λ-1=1.097*107(3-2-5-2)m-1

λ-1=7.8009*105m-1

λ=(7.8009*105m-1)-1

λ=1.282*10-6m